# SDEval: Safety Dynamic Evaluation for Multimodal Large Language Models

### v1

@@ 4.2 Results of Dynamic Evaluation
To understand the impact of each strategy, we conduct comprehensive experiments to figure out the most powerful dynamic operations in Section 4.3. We select the most influential strategies (Word Replacement and Figstep) to conduct experiments on MLLMGuard (Gu et al. 2024) and VLSBench (Hu et al. 2024).
@@
| Strategy | Variants | ASD() | PAR() | ||
| Oringinal | Vanilla | 32.21 | 40.19 | ||
Word Replacement | 38.71 (6.300) | 26.94 (13.25) | ||
| Sentence Paraphrasing | 36.68 (4.270) | 26.07 (14.12) | |||
| Adding Descriptions | 32.97 (0.560) | 30.67 (9.520) | |||
| Making Typos | 35.56 (3.150) | 30.70 (9.490) | |||
| Linguistic Mix | 33.23 (0.820) | 32.95 (7.240) | |||
| Chain-of-Thought | 32.67 (0.260) | 30.96 (9.230) | |||
Adding Texts | 33.73 (1.320) | 29.58 (10.61) | ||
| Adding Objects | 39.41 (7.000) | 26.45(13.74) | |||
| Generation | 34.89 (2.480) | 26.14 (14.05) | |||
| Augmentation | 33.97 (1.560) | 36.46 (3.730) | |||
| Style Transfer | 35.20 (2.790) | 26.38 (13.81) | |||
Text-to-Image | 35.10 (2.690) | 24.36 (15.83) | ||
| Image-to-Text | 34.89 (2.480) | 31.65 (8.540) | |||
| FigStep | 41.96 (9.550) | 17.08 (23.11) | |||
@@
As shown in Table 2, the closed-source model performs much better than open-source models in both original and dynamic benchmarking. Claude-4-Sonnet (Anthropic. 2025) gets the highest safety rate, and it is also safer than others. Additionally, after the dynamic strategies, the safety of all MLLMs decreases, and the proportion of answers judged as containing warnings decreases more, which shows that the dynamic strategy causes more safety risks for MLLMs, thereby reducing the safety rate of the model. The dynamic evaluation results show that the MLLMs still face huge safety risks.
@@
Overall, these results show that current MLLMs are not good enough to cope with safety dynamic evaluation, suggesting there is data leakage in the current model training process, and current MLLMs still can’t handle safety issues well. How to ensure that the model’s safety and performance can develop in a balanced manner under the AI Law (Yang et al. 2024a) roadmap is still a huge challenge.
@@ 4.3 Ablation Study
In order to explore the impact of each dynamic strategy on safety evaluation, we select some powerful open-source multimodal large language models, including Qwen-VL family (Bai et al. 2025), InternVL family (Zhu et al. 2025), Yi-VL family (Young et al. 2024), to conduct comprehensive ablation experiments for each dynamic operation based on MLLMGuard (Gu et al. 2024) Benchmark. We only show the results of InternVL-Chat-V1.5 in Table 3. More ablation experiment results can be found in our Appendix.
@@ 5 Safety-Capability Balance
As SDEval is a general framework, it can also be utilized for MLLM capability evaluation. The input for the capability evaluation of MLLM still consists of two parts: text and images. We can utilize the proposed text, image, and text-image dynamic strategies to conduct capability evaluation for MLLM, which can relieve the data leakage issues. And here we choose the same dynamic strategies as safety dynamic evaluation for capability evaluation. We apply SDEval dynamic evaluation on some popular benchmark datasets (MMVet (Yu et al. 2023), MMBench (Liu et al. 2023d)) to assess current MLLMs. As can be seen in Table 4, after applying the proposed dynamic strategies, all the metrics on MMVet and MMBench decreased, which means that our SDEval framework is not only suited for safety evaluation, but also suited for more general evaluation benchmarks. More results can be seen in our Appendix.
@@
AI theory (Zhang et al. 2025b; Yang et al. 2024a) hypothesizes that the development of AGI should consider the balance of model performance and safety. The safety and capability of AI are generally balanced along a roadway. In the short term, rotation is allowed, but in the long term, it should not be lower than , as in the current state, or higher than , which would hinder development and industrial application. We weighted the safety and capability scores of the selected datasets after dynamic evaluation according to the dataset size and drew a capability-safety scatter plot based on this. As can be seen in Figure 5 (a), we present the weighted ranking figure: Claude-4-Sonnet outperforms all the models on safety, and it also has a good performance in the intelligent capability, and Gemini-2.5-Pro has achieved an excellent balance between safety and capability. As presented in Figure 5 (b), most MLLMs have worse robustness in safety, resulting in more significant safety performance loss when performing dynamic strategies, which highlights the need to strengthen the model’s safety ability in future development.

### v2

@@ 4.2 Results of Dynamic Evaluation
To understand the impact of each strategy, we conduct comprehensive experiments to figure out the most powerful dynamic operations in Section 4.3. We select the most influential strategies (Word Replacement and Figstep) to conduct experiments on MLLMGuard (gu2024mllmguard) and VLSBench (hu2024vlsbench).
@@
| Strategy | Variants | ASD() | PAR() |
| Oringinal | Vanilla | 32.21 | 40.19 |
| Text Dynamics | Word Replacement | 38.71 (6.300) | 26.94 (13.25) |
| Sentence Paraphrasing | 36.68 (4.270) | 26.07 (14.12) | |
| Adding Descriptions | 32.97 (0.560) | 30.67 (9.520) | |
| Making Typos | 35.56 (3.150) | 30.70 (9.490) | |
| Linguistic Mix | 33.23 (0.820) | 32.95 (7.240) | |
| Chain-of-Thought | 32.67 (0.260) | 30.96 (9.230) | |
| Image Dynamics | Adding Texts | 33.73 (1.320) | 29.58 (10.61) |
| Adding Objects | 39.41 (7.000) | 26.45(13.74) | |
| Generation | 34.89 (2.480) | 26.14 (14.05) | |
| Augmentation | 33.97 (1.560) | 36.46 (3.730) | |
| Style Transfer | 35.20 (2.790) | 26.38 (13.81) | |
| Text-Image Dynamics | Text-to-Image | 35.10 (2.690) | 24.36 (15.83) |
| Image-to-Text | 34.89 (2.480) | 31.65 (8.540) | |
| FigStep | 41.96 (9.550) | 17.08 (23.11) | |
@@
As shown in Table 2, the closed-source model performs much better than open-source models in both original and dynamic benchmarking. Claude-4-Sonnet (claude_4_opus) gets the highest safety rate, and it is also safer than others. Additionally, after the dynamic strategies, the safety of all MLLMs decreases, and the proportion of answers judged as containing warnings decreases more, which shows that the dynamic strategy causes more safety risks for MLLMs, thereby reducing the safety rate of the model. The dynamic evaluation results show that the MLLMs still face huge safety risks.
@@
Overall, these results show that current MLLMs are not good enough to cope with safety dynamic evaluation, suggesting there is data leakage in the current model training process, and current MLLMs still can’t handle safety issues well. How to ensure that the model’s safety and performance can develop in a balanced manner under the AI Law (yang2024towards) roadmap is still a huge challenge.
@@ 4.3 Ablation Study
In order to explore the impact of each dynamic strategy on safety evaluation, we select some powerful open-source multimodal large language models, including Qwen-VL family (qwen25vl), InternVL family (internvl3), Yi-VL family (Young2024YiOF), to conduct comprehensive ablation experiments for each dynamic operation based on MLLMGuard (gu2024mllmguard) Benchmark. We only show the results of InternVL-Chat-V1.5 in Table 3. More ablation experiment results can be found in our Appendix.
@@ 5 Safety-Capability Balance
As SDEval is a general framework, it can also be utilized for MLLM capability evaluation. The input for the capability evaluation of MLLM still consists of two parts: text and images. We can utilize the proposed text, image, and text-image dynamic strategies to conduct capability evaluation for MLLM, which can relieve the data leakage issues. And here we choose the same dynamic strategies as safety dynamic evaluation for capability evaluation. We apply SDEval dynamic evaluation on some popular benchmark datasets (MMVet (yu2023mm), MMBench (Liu2023MMBenchIY)) to assess current MLLMs. As can be seen in Table 4, after applying the proposed dynamic strategies, all the metrics on MMVet and MMBench decreased, which means that our SDEval framework is not only suited for safety evaluation, but also suited for more general evaluation benchmarks. More results can be seen in our Appendix.
@@
AI theory (AIBench; yang2024towards; zhang2025xiongkuo) hypothesizes that the development of AGI should consider the balance of model performance and safety. The safety and capability of AI are generally balanced along a roadway. In the short term, rotation is allowed, but in the long term, it should not be lower than , as in the current state, or higher than , which would hinder development and industrial application. We weighted the safety and capability scores of the selected datasets after dynamic evaluation according to the dataset size and drew a capability-safety scatter plot based on this. As can be seen in Figure 5 (a), we present the weighted ranking figure: Claude-4-Sonnet outperforms all the models on safety, and it also has a good performance in the intelligent capability, and Gemini-2.5-Pro has achieved an excellent balance between safety and capability. As presented in Figure 5 (b), most MLLMs have worse robustness in safety, resulting in more significant safety performance loss when performing dynamic strategies, which highlights the need to strengthen the model’s safety ability in future development.